OpenStudy (anonymous):

Which values from below aren't in the doamin of f(x)=((x^2+2x)/(x^2-9)) -1,-3,-2,0,1,2,3

hartnn (hartnn):

the values which make the denominator =0 will not be in the domain.

hartnn (hartnn):

so, solve for x^2-9=0

OpenStudy (anonymous):

3

hartnn (hartnn):

and ?

OpenStudy (anonymous):

0

hartnn (hartnn):

how 0 ?

hartnn (hartnn):

0^2-9 will not be =0

OpenStudy (anonymous):

sorry 3

OpenStudy (anonymous):

hit the wrong number

hartnn (hartnn):

x^2-9 = 0 (x+3)(x-3)=0 x+3 =0 , x-3=0 x=..., x=.....

hartnn (hartnn):

one of the number is 3 is correct, oter ?

OpenStudy (anonymous):

x=3 and x=-3

hartnn (hartnn):

now its correct :)

OpenStudy (anonymous):

thank you

hartnn (hartnn):

welcome ^_^